5.At the instant, traffic light turns green, an automobile that has been waiting at an intersection of roads accelerate with 5m/s. At the same instant a truck travelling with a constant velocity of 20m/s overtake and passes the automobile.
a.How far beyond it's starting point will the automobile overtake the truck.
b.How fast will it be travelling?
1
Expert's answer
2021-08-11T09:53:45-0400

"\text{for automobile:}"

"v_0 =0;a= 5\frac{m}{s^2}"

"s_{aut}(t) = v_0t+\frac{at^2}{2}=\frac{5t^2}{2}"

"v_{aut}(t)= v_0+at = 5t"

"\text{for truck:}"

"v_{tr}= 20\frac{m}{s}"

"s_{tr}(t) = v_{tr}t= 20 t"


"s_{aut} = s_{tr}"

"\frac{5t^2}{2}= 20t"

"t_1= 0 \text{ this is the beginning of the movement}"

"t_2 = 8s"

"s_{tr}(t_2) = 20 *t_2= 160m"

"v_{aut}(t_2)=5t_2= 40\frac{m}{s}"


"\text{Answer: }a."160 m b."40\frac{m}{s}"